in a class the number of boys to girls is in the ratio of 7:8.if they were increased by 20 and 10 percentage .what would be the new  ratio?
 

Grade:9

1 Answers

Ravi
9 years ago
Boys= 7x
Girls= 8x

New strength of boys=7x(1+20/100)=42x/5
New strength of girls=8x(1+10/100)=44x/5

So new ratio will be (42x/5)(44x/5)
=(42/44)=21/22
....................................................................................................................................